Rick's cafe ordered coffee beans from two different distributors at different prices. Rick wants to combine the two types of beans to make a mixture that costs $2.02 per pound. How many pounds of coffee beans that cost $3.00 per pound should be mixed with 7 pounds of coffee beans that cost $1.60 per pound to obtain the desired mixture?
ok, just a big old brain lapse. Thanks.

Answer by Paul(988)

300x+160(7-x)=202(7)
300x+1120-160x=1414
140x=294
x=2.1
7-2.1=4.9
hence for 3.00 it should be 2.1 and for 1.60 it should be 4.9